OpenStudy (miah0420):

Look at the two triangular flags below. What is the length of the side BC of triangle ABC?

OpenStudy (anonymous):

no triangle yet

OpenStudy (anonymous):

5x - 2 = x + 6 4x = 8 x = 2 5*2 - 2 = 10-2= 8cm

OpenStudy (anonymous):

dont forget my medal

OpenStudy (anonymous):

what grade is that by the way